OpenStudy (anonymous):

Solve 5(x – 5) = 4(x – 1) + 3x.

OpenStudy (uri):

5(x – 5) = 4(x – 1) + 3x. 5x-25=4x-4+3x 5x-4x-3x=-4+25 -3x=21 x=21/-3 x=-7

OpenStudy (anonymous):

5(x – 5) = 4(x – 1) + 3x 5x-25=4x-4+3x 5x-3x-4x=-4+25 -2x=21 x=-10.5

OpenStudy (anonymous):

uri is wrong

OpenStudy (uri):

Ah it's -2x* not -3x,Sorry :p